OpenStudy (anonymous):

Solve the equation using the Zero-Product Property -9n(5n-5)=0 a. -1/9, 1 b.0, 1 c. -1/9, -1 d. 0, -1

OpenStudy (johnweldon1993):

right, the zero product property.. in your equation -9n(5n-5) = 0 there are 2 ways to make this = 0 correct? how can -9n = 0? if 'n' = what number? and how can 5n - 5 = 0? if 'n' = what number?

OpenStudy (campbell_st):

well then you need to solve 9n = 0 and 5n - 5 = 0 solve both equations for n